OpenStudy (anonymous):

a/2 - a/5 + 7= 10, a=? Heelp me.

OpenStudy (anonymous):

fractions are annoying, the way we deal with more than one fraction is we make a common denominator, which means we try to get the bottom number the same on each fraction for example \[\frac{1}{2} + \frac{1}{3} = \frac{3}{6} + \frac{2}{6} = \frac{5}{6}\] in this case we have a/2 and a/5, so we want to get them both over the lowest common denominator which is 10 \[\frac{a}{2} = \frac{5a}{10}\] can you do the other?